# appAppendix References

### v5

@@ 3.3 Time-dependent Classifier
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process. Assume we condition on the label $y=c$. Firstly, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack. Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set. For each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ denotes the cross-entropy loss) and use the gradient and the target DPM to reverse the diffusion process. Next, we compute the similarity score for each generated image. Finally, we evaluate the attack performance using evaluation metrics. Defer to Appendix E for SIDE’s pseudocode.
@@ 4 Experiments
In this section, we first describe our experimental setup, introduce the performance metrics and then present the main evaluation results of our SIDE method. We empirically verify that the memorized images are not from the classifier. We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
@@ 4.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [23] which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) [21] which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer [22] with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM [33]) using the HuggingFace implementation [36]. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 [15] as the normal classifier.
4.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection [31, 13]. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
The AMS and UMS metrics are defined as follows:
$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{1},\alpha,\beta\right)}}{N_{G}}$ (5)
$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}}$ (6)
where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent:
@@ 4.2 Performance Metrics
where $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following [5], while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. (examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ ditance [5] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
@@
The AMS averages the similarity scores across generated images, ensuring that memorized images are not overlooked. In contrast, the UMS quantifies distinct memorized instances by evaluating unique matches, thereby accounting for the uniqueness of the memorized images. By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.

### v6

@@ 3.4 Overall Procedure of SIDE
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE method extracts training data from the model through a conditional generation process, as described in Algorithm 1. Given the label $y=c$, the complete procedure of SIDE is as follows. It begins by selecting a set of values for $\lambda$, denoted as $\mathcal{S}_{\lambda}$ (Line 1). For each $\lambda$, the target DPM is required to generate $N_{G}$ data samples (Lines 18 - 27). Then, at each sampling timestep $t$, the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ (Line 22) is computed, where $CE(\cdot)$ is the cross-entropy loss. This gradient, together with the target DPM, is then used to guide the reverse diffusion process. For each generated image, a similarity score is computed against every ground-truth training images to identify the most closely matching original image.
Using the similarity scores obtained for the extracted images, a performance metric can then be calculated to evaluate the effectiveness of the extraction process. Existing methods predominantly rely on $L_{p}$ distance metrics to identify pixel-wise similar samples. However, such metrics are insufficient for capturing semantically meaningful similarities. In the experiment section, we will introduce a set of performance metrics to facilitate a more comprehensive evaluation of the extraction methods.
Algorithm 1 SIDE
1:Input: DPM $p_{\theta}(\bm{x})$, classifier $p_{\theta}(y_{I}|\bm{x})$, $\mathcal{S}_{\lambda}$, $N_{G}$, label $c$, $N_{syn}$, $T$, Denoise Step $DS(\bm{x}_{t},t,\bm{s})$
2:Output: Extracted data $\mathcal{D}_{ext}$, classifier $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
3:Classifier Training:
4:Step 1: Architecture
5:Modify $p_{\theta}(y|\bm{x})$ to $p_{\theta}^{t}(y|\bm{x}_{t})$ for time input
6:Step 2: Distillation
7:for $i=1$ to $N_{syn}$ do $\triangleright$ Generate synthetic data
8:    $\bm{x}_{T}\sim\mathcal{N}(0,I)$
9:    for $t=T$ down to $1$ do
10:         $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\theta}(\bm{x}_{t},t))$
11:    end for
12:    $y_{I}^{(i)}\leftarrow p_{\theta}(y_{I}|\bm{x}_{0}^{(i)})$triangleright Generate pseudo-labels
13:end for
14:Step 3: Train $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$:
15:$\mathcal{L}_{distill}=\frac{1}{N_{syn}}\sum_{i=1}^{N_{syn}}D_{KL}(p_{\theta}(y_{I}|\bm{x}_{0}^{(i)}),p_{\theta}^{t}(y_{I}|\bm{x}_{t}^{(i)}))$
16:Data Extraction:
17:Initialize $\mathcal{D}_{ext}\leftarrow\emptyset$
18:for $\lambda\in\mathcal{S}_{\lambda}$ do
19:    for $i=1$ to $N_{G}$ do
20:         $\bm{x}_{T}\sim\mathcal{N}(0,I)$
21:         for $t=T$ down to $1$ do
22:             $g_{t}\leftarrow\nabla_{\bm{x}_{t}}CE(c,p_{\theta}^{t}(y_{I}|\bm{x}_{t}))$
23:             $s_{\text{guided}}\leftarrow\nabla_{\bm{x}_{t}}\log p_{\theta}^{t}(\bm{x}_{t})-\lambda g_{t}$
24:             $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\text{guided}})$
25:         end for
26:         Append $\bm{x}_{0}$ to $\mathcal{D}_{ext}$
27:    end for
28:end for
29:return $\mathcal{D}_{ext}$, $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
4 Theoretical Analysis
In this section, we first introduce a Kullback-Leibler (KL) divergence-based distribution measure to quantify the degree of memorization in generative models. Building on this measure, we provide a theoretical explanation for data memorization in conditional diffusion probabilistic models (DPMs) and elucidate why SIDE is effective for unconditional DPMs.
4.1 Distributional Memorization Measure
There are various approaches to measure the memorization effect of a generative model. One approach involves comparing each generated sample with the raw training samples in a sample-by-sample manner, as previously mentioned. While such measures (e.g., $L_{p}$ distances) are effective for evaluating data extraction performance, they are insufficient for analyzing the overall memorization effect of the model. To characterize the model-level memorization effect with respect to the training data distribution and to support our theoretical analysis, we introduce the following distributional memorization measure.
To assess the model’s memorization ability with respect to the overall data distribution while also accounting for the memorization of each individual data point, it is natural to consider the Kullback-Leibler (KL) divergence between the uniform empirical distribution over the dataset $\mathcal{D}$, defined as $\frac{1}{|\mathcal{D}|}\sum_{x_{i}\in\mathcal{D}}\delta(x_{i})$ with $\delta(\cdot)$ is the Dirac delta function, and the distribution $p$ induced by the model’s generated samples. Here, the $\delta(\cdot)$ function enforces a point-wise memorization measure, which quantifies how closely the generated samples align with each individual data point in the original distribution. A smaller KL divergence indicates that the model exhibits stronger memorization of the training data points. However, for a continuous probability distribution $p$, computing such a KL divergence directly is infeasible. To address this, we approximate each Dirac delta function by a normal distribution with a small variance in the following definition.
Definition 1 (Memorization Divergence).
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset $\mathcal{D}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of divergence between $p_{\theta}$ and distribution of training dataset is defined as:
$\mathcal{M}(\mathcal{D};p_{\theta},\epsilon)$ $=D_{\mathrm{KL}}(q_{\epsilon}\|p_{\theta})$ (5)
with $\text{with }q_{\epsilon}(x)=$ $\frac{1}{N}\sum_{x_{i}\in\mathcal{D}}\mathcal{N}(x|x_{i},\epsilon^{2}I),$
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ is the normal distribution with mean ${\bm{x}}_{i}$ and covariance matrix $\epsilon^{2}I$.
Note that in Eq. (5), a smaller value of ${\mathcal{M}}({\mathcal{D}};p_{\theta},\epsilon)$ indicates greater overlap between the two distributions, reflecting stronger memorization. As $\epsilon$ approaches 00, the measured memorization divergence becomes more accurate. In fact, the normal distribution $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ can be replaced by any family of continuous distributions $\hat{q}(x|x_{j},\epsilon)$ that satisfies two conditions: 1) symmetric with respect to $x$ and $x_{j}$(i.e., $\hat{q}(x|x_{j},\epsilon)=\hat{q}(x_{j}|x,\epsilon)$) and 2) convergence to $\delta(x_{i})$ in distribution. This substitution does not affect the result of Theorem 1 in Sec. 4.2.
While concerns may arise regarding the influence of $\epsilon$ on the divergence, this divergence is primarily intended for comparative analysis when $\epsilon$ is sufficiently small. When comparing the memorization divergence across different models, the $\epsilon$ will not influence relative differences, as demonstrated in Appendix A.
4.2 Theoretical Analysis
Building on the memorization divergence measure, here we provide a theoretical analysis to explain why conditional DPMs exhibit stronger memorization effect. Our analysis centers on the concept of informative labels, which represent information capable of partitioning a dataset into multiple disjoint subsets. We show that DPMs, conditioned on informative labels, exhibit stronger memorization effects.
Informative Labels
The concept of informative labels has been previously discussed as class labels [18]. In this work, we extend this notion by introducing a more general definition that encompasses includes both class labels and random labels as special cases. Formally, we define an informative label as follows.
Definition 2 (Informative Label).
Let ${\mathcal{Y}}$ be an attribute of the data that takes values in $\{y_{i}\}_{i=1}^{C}$. We define ${\mathcal{Y}}$ as an informative label if it enables the partitioning of the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$, where each subset corresponds to a distinct value of ${\mathcal{Y}}$.
In the above definition, informative labels are not restricted to conventional class labels; they can also include text captions, features, or cluster information that can be used to group training samples into subsets. The definition emphasizes that an informative label must be capable of distinguishing a subset of samples from others. An extreme case occurs when all samples share the same label, rendering the label non-informative. According to our definition, both class-wise and random labels are special cases of informative labels. Informative labels can be either explicit, such as class labels, random labels, or text captions, or implicit, such as salient features or clusters.
Next, we present our main theoretical result on the memorization mechanism of conditional DPMs and explain why SIDE is effective. Let ${\mathcal{D}}_{i}$ denote the subset consisting of data with the informative label ${\mathcal{Y}}=y_{i}$. Additionally, we denote the data distribution in the original dataset ${\mathcal{D}}$ by $p$. Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the data distribution corresponding to that attribute.
Theorem 1.
If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ almost everywhere for the informative label $y_{i}$, i.e., $TV(p_{i},p_{\theta_{i}})=0$, then with probability 1:
$\lim_{\epsilon\to 0}\lim_{|{\mathcal{D}}_{i}|\to\infty}\left({\mathcal{M}}({\mathcal{D}}_{i};p_{\theta_{i}},\epsilon)-{\mathcal{M}}({\mathcal{D}}_{i};p_{\theta},\epsilon)\right)\leq 0,$ (6)
where $TV(\cdot)$ is the total variance distance, $p_{\theta_{i}}$ and $p_{\theta}$ denote the distribution of generated data for model trained on data labeled $y_{i}$ and the whole dataset. The equality holds if and only if $TV(p,p_{i})=0$.
The proof for Theorem 1 is in Appendix A. The theorem implies that any form of informative labels, including class labels and random labels, can enhance the memorization effect of conditional generative models. Conditional models exhibit stronger memorization by isolating data subsets through explicit informative labels, allowing focused training on each group’s unique patterns, which sharpens their ability to memorize specific training examples. SIDE succeeds by uncovering and exploiting the unconditional DPM’s inherent but implicit informative labels, using classifier guidance to concentrate generation on specific memorized features.
Table 1: The AMS (%) and UMS (%) results at low, mid, and high similarity levels. ‘Random’ is the baseline method proposed in [7], while OL-TI is SIDE without TDKD training. We also compare SIDE using classifiers trained on the original (SIDE-O) vs. disjoint auxiliary (SIDE-A) datasets. The similar performance between SIDE-O and SIDE-A suggests there is no training data leaking via the classifier.
Dataset Method Low Similarity Mid Similarity High Similarity
AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
SIDE-A 5.250 2.020 2.300 0.880 1.620 0.640
SIDE-O 5.320 2.050 2.490 0.860 1.770 0.560
CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
SIDE-A 15.010 2.624 1.310 0.554 0.090 0.082
SIDE-O 15.172 2.342 1.115 0.444 0.054 0.044
CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
SIDE-A 8.712 6.802 0.234 0.234 0.010 0.010
SIDE-O 8.756 6.940 0.224 0.212 0.012 0.012
CelebA Random 1.953 1.895 0.000 0.000 0.000 0.000
OL-TI 1.145 1.125 0.000 0.000 0.000 0.000
SIDE-A 4.682 4.706 0.098 0.098 0.000 0.000
SIDE-O 4.883 4.766 0.078 0.078 0.000 0.000
ImageNet Random 0.000 0.000 0.000 0.000 0.000 0.000
OL-TI 0.000 0.000 0.000 0.000 0.000 0.000
SIDE-A 0.152 0.152 0.076 0.076 0.000 0.000
SIDE-O 0.114 0.114 0.038 0.038 0.000 0.000
Figure 3: A comparison between the original training images (top row) and extracted images (bottom row) by our SIDE method. The matches are classified into three categories: low similarity (SSCD score $<$ 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score $>$ 0.6). This highlights varying degrees of semantic resemblance among the image pairs.
@@ 5 Experiments
In this section, we first introduce the experimental setup and performance metrics, followed by the main evaluation results. We also show that the memorized images are not derived from the classifier. Additionally, we conduct an ablation study and hyperparameter analysis to provide deeper insights into the working mechanism of SIDE.
@@ 5.1 Experimental Setup
We evaluate the effectiveness of our method on five datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [30], which consists of 5478 images with a resolution of 128$\times$128; 2) a subset of the CelebA (CelebA-25000) [28], containing 25,000 images with a resolution of 128$\times$128; 3) CIFAR-10, containing 50,000 images with a resolution of 32$\times$32; 4) CelebA [28], which contains 22059 images with a resolution of 128$\times$128; and 5) ImageNet [14], which contains $\approx$ 14 million images with a resolution of 256$\times$256. All the images are normalized to the range of [-1,1]. We use the AdamW optimizer [29] with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM [41]) using the HuggingFace implementation [44]. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and CelebA, 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, and 1980K steps on ImageNet, respectively. We use ResNet34 [20] as the pseudo-labeler. For ImageNet, the DPM is trained on the entire dataset. However, due to the scale of ImageNet, we only evaluate on a smaller subset known as ImageNette [23].
5.2 Image-level Performance Metrics
As previously mentioned, determining whether an extracted image is a memorized copy of a training image is challenging, and $L_{p}$ distances in pixel space are ineffective when the image is semantically similar but not an exact copy. Prior research [39, 18] addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. To address these issues, we propose two new memorization scores: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS). The AMS and UMS scores are defined as follows:
$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)}}{N_{G}}$ (7)
$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}},$ (8)
where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, they can be formulated as:
@@ 5.2 Image-level Performance Metrics
where $\gamma$ denotes the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$, following [7], while for high-resolution datasets, we use the SSCD score as $\gamma$.
@@
The AMS averages the similarity scores across extracted images, ensuring that memorized images are not overlooked. In contrast, the UMS quantifies distinct memorized images by evaluating unique matches, thereby accounting for the uniqueness of the memorized images. By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. (examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ distance [7] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.